[ How many hours will a car traveling at 40 miles per hour be driven before being overtaken by another car leaving from the same place one hour later traveling at 50 miles per hour? ]
I also need a formula for this problem.

Answer by TimothyLamb(4379)

s = d / t
d = s * t
---
when they meet, both cars are at the same distance (D) from the departure point
D = 40t
D = 50(t - 1)
D = 50t - 50
---
D = 40t = 50t - 50
10t = 50
---
t = 5 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php